jigglypuff314 (jigglypuff314):

\[\int\limits_{0}^{1}xe^{-x}~dx = \]a) 1-2e b) -1 c) 1-2e^(-1) d) 1 e) 2e-1

jigglypuff314 (jigglypuff314):

Could someone check my work pwease? :3 u = x v = -e^(-x) du = dx dv = e^(-x) dx uv - ∫v du -xe^(-x) + ∫e^(-x) dx -xe^(-x) - e^(-x) -e^(-1) - e^(-1) - (-e^(0)) -2e^(-1) + 1 so C ?

OpenStudy (anonymous):

\[I=\int\limits_{0}^{1}x~e ^{-x}dx=[ x \frac{ e ^{-x} }{ -1 }-\int\limits_{0}^{1} 1~\frac{ e ^{-x} }{ -1 }dx ]from~0\rightarrow1\] \[=\left[ -\frac{ x }{ e^x }+\frac{ e ^{-x} }{ -1 } \right]~from~0\rightarrow1\] \[=-\left[ \frac{ x+1 }{ e^x } \right]~from~0\rightarrow1\] \[=-\left[ \frac{ 1+1 }{ e^1 }-\frac{ 0+1 }{ e^0 } \right]=-\frac{ 2 }{ e }+1=1-2 e ^{-1}\] so you are correct.

jigglypuff314 (jigglypuff314):

Thank you! :)

OpenStudy (anonymous):

yw